# False Sense of Security in Explainable Artificial Intelligence (XAI)

### v1

@@ 4.0.4 Data and Concept Drift
Fourth, data drift Maimon and Rokach (2010) and concept drift Widmer and Kubat (1996) refer to the change in data distribution and in applicable domains, respectively. Once trained and deployed, AI systems with explainers tend to remain fixed. Nonetheless, data and their relationship to underlying concepts fed to the systems may change over time (Figure 4). For example, an AI model trained with data from African hospitals may end up being used in Asian countries. Patients of a different ethnicity impact how AI is performing, as well as what explanations are given. XAI that worked faithfully in Africa might gives misleading explanations in Asia. On the other hand, over time, new concepts may appear, which originally did not exist in the training data (Figure 4(d)). For example, malignant tumors among Asian patients might exhibit novel characteristics in medical images, unseen among African patients. Then, XAI methods that trained and evaluated on African populations might not work on medical images from Asian patients. To overcome these challenges, we must evaluate XAI for application under consideration and continuously update our methods over time. Continuous data monitoring, collection, training, evaluation and deployment would involve significant costs and effort.
@@ 5 Market Conditions
Note that over-confident marketing of XAI are not at all unique to these two companies. Explainability methods have been developed in a research environment, often within narrowly defined applications and assumptions. As those methods are packaged in a for-profit environment, technical nuances may have been glanced over.
@@ 6 Discussion and Conclusion
Regulators and politicians are under the impression that explainability is an issue of implementation, instead of a fundamental research problem. Law and policy documents fail to specify what are and how to obtain explanations, while simultaneously requiring certain AI operators to provide explanations. The AI Act in EU European Commission (2021) and the Executive Order in the USA Biden (2023) demands safe, truthworthy, and explainable AI. These will force tech companies to implement and outsource explainability for their AI systems, regardless of their accuracy, reliability, or utility.
@@
Clients and citizens are conditioned to rely and trust the AI systems on the basis of sweeping laws and AI systems promising explanations. Whereas a black box AI raises an inherent concern about their unknowable operating characteristics, legally mandated and confidently marketed XAI, is provided with a false aura of trustworthiness. User studies show inclusion of explanations increases user trust and decrease decision making time. But at the same time, XAI (specifically, LIME Ribeiro et al. (2016), SHAP Lundberg and Lee (2017), and TreeInterpreter Saabas (2014)) substantially decreased the decision accuracy of human analysts in a fraud detection study Jesus et al. (2021). Regardless of the model performance, users of AI becomes more reliant when explanations are provided de Brito Duarte et al. (2023).
From both technical and legislative perspectives, explainability plays an important role for AI governance. While it is intuitive to present and advocate for the need to understand, interpret, and explain a model’s decision making process, the methodologies of XAI require thorough consideration of their applications, goals, and failure modes. We must dispel the false sense of security in XAI arising from over-confident products that seemingly satisfy legislative and policy texts that vaguely imply a right to explanation.

### v2

@@ 4.0.4 Data and Concept Drift
Fourth, data drift Maimon and Rokach (2010) and concept drift Widmer and Kubat (1996) refer to the change in data distribution and in applicable domains, respectively. Once trained and deployed, AI systems with explainers tend to remain fixed. Nonetheless, data and their relationship to underlying concepts fed to the systems may change over time (Figure 4). For example, an AI model trained with data from African hospitals may end up being used in Asian countries. Patients of a different ethnicity impact how AI is performing, as well as what explanations are given. XAI that worked faithfully in Africa might gives misleading explanations in Asia. On the other hand, over time, new concepts may appear, that originally did not exist in the training data (Figure 4(d)). For example, malignant tumors among Asian patients might exhibit novel characteristics in medical images, unseen among African patients. Then, XAI methods that were trained and evaluated on African populations might not work on medical images from Asian patients. To overcome these challenges, we must evaluate XAI for the application under consideration and continuously update our methods over time. Continuous data monitoring, collection, training, evaluation and deployment would involve significant costs and effort.
@@ 5 Market Conditions
Note that over-confident marketing of XAI is not at all unique to these two companies. Explainability methods have been developed in a research environment, often within narrowly defined applications and assumptions. As those methods are packaged in a for-profit environment, technical nuances may have been glanced over.
@@ 6 Discussion and Conclusion
Regulators and politicians are under the impression that explainability is an issue of implementation, instead of a fundamental research problem. Law and policy documents fail to specify what are and how to obtain explanations, while simultaneously requiring certain AI operators to provide explanations. The AI Act in EU European Commission (2021) and the Executive Order in the USA Biden (2023) demand safe, trustworthy, and explainable AI. These will force tech companies to implement and outsource explainability for their AI systems, regardless of their accuracy, reliability, or utility.
@@
Clients and citizens are conditioned to rely and trust the AI systems on the basis of sweeping laws and AI systems promising explanations. Whereas a black box AI raises an inherent concern about their unknowable operating characteristics, legally mandated and confidently marketed XAI, is provided with a false aura of trustworthiness. User studies show inclusion of explanations increases user trust and decreases decision-making time. But at the same time, XAI (specifically, LIME Ribeiro et al. (2016), SHAP Lundberg and Lee (2017), and TreeInterpreter Saabas (2014)) substantially decreased the decision accuracy of human analysts in a fraud detection study Jesus et al. (2021). Regardless of the model performance, users of AI become more reliant when explanations are provided de Brito Duarte et al. (2023).
From both technical and legislative perspectives, explainability plays an important role in AI governance. While it is intuitive to present and advocate for the need to understand, interpret, and explain a model’s decision-making process, the methodologies of XAI require thorough consideration of their applications, goals, and failure modes. We must dispel the false sense of security in XAI arising from over-confident products that seemingly satisfy legislative and policy texts that vaguely imply a right to explanation.